# When Large Language Models Meet Vector Databases: A Survey

### v1

@@ 3.1 Vector Databases, the V-factor
Traditional full-text search engines, such as Elastic Search and Amazon OpenSearch, are based on term frequency metrics like BM25, which has proved practical for many search applications. However, such search techniques require significant investment in time and expertise to tune them to account for the meaning or relevance of the terms searched. On the other hand, vector databases successfully solved the aforementioned problem, thus the emergence of vector databases (as shown in Table 3) has greatly influenced machine learning systems and their multi-modal applications. An intuitive application using vector search is searching images with an image on search engines built on vector databases Wang and et al. (2021) like Google Image Search  [^2] https://cloud.google.com/vertex-ai/docs/vector-search/overview, which uses one input image to find similar images on the internet. While vector databases boast their unique search capability and efficiency for unstructured data, they are just a backend factor for manifold applications. To maximize their abilities, this leads us to an interesting question, since LLMs use embeddings to represent text as vectors, can developers combine vector databases and LLMs to overcome the aforementioned challenges that inherit in LLM applications? The answer is yes.
@@
As a kind of database encapsulated with vector search in vector data that represents real-world information within high dimensionalities, vector databases are well-capable for retrieval applications Asai and et al. (2023) incorporating LLMs because LLM applications are generally read-intensive, not requiring many write-related changes, especially data deletes Pan et al. (2023). On the other hand, vector databases can efficiently manage and warehouse vector data required and generated by LLMs, thus providing a solid data cornerstone for both LLMs and their applications. While LLMs are often limited by domain knowledge that cannot be uploaded or distributed due to security and privacy concerns, domain knowledge is often unstructured data that can easily embedded into vector databases for efficient local retrieval and further integration with generative AIs. Moreover, the computing and storage resources of vector databases are way cheaper than LLMs, since it does require costly GPUs and TPUs, thus achieving a cost-effective way of fast retrieval and durable (non-volatile) storage.
3.2 Retrieval-Augmented Generation (RAG): VecDB as an External Knowledge Base
@@ 3.2.1 Development of RAG Paradigm
As the release of ChatGPT casts spotlight on LLMs to the public world, there is an increasing need of using AI chatbots as query and retrieval agents. But simply loading users’ private data as input to LLMs is found to be incompetent in real-world applications. Because LLMs have been constrained by their limited token counts and the high costs associated with training and fine-tuning for every alterations of data, especially pronounced when dealing with personalized or business-specific responses that require real-time data updates. To address such concerns and need, retrieval-augmented generation (RAG) emerges as a novel solution that addresses the challenges faced by LLMs in integrating and processing large and dynamic data in external databases, where vector databases offer a solution by acting as external memory for LLMs. They allow for the segmentation of private data by converting them into vectors and storing these vectors efficiently for quick retrieval process. Integrating with vector databases enables LLMs to access and synthesize enormous amount of data without the need for constant re-training, thereby overcoming their inherent limitations.
@@
The concept of RAG is devised to be a paradigm, and a common workflow of RAG is illustrated in Figure 1. There are essentially 3 main parts to complete one full run of the systems: data storage, retrieval, and generation.
@@
Data storage means establishing reliable external memory like vector databases, and there are detailed recipes for this process Han et al. (2023). It starts with the data preprocessing, during which the original data is collected, cleaned, integrated, transformed, normalized and standardized. The processed data is chunked into smaller segments because of the contextual limit of LLMs. These segments of data are then converted by an embedding model into vectors, the semantic representations of the data, which are stored in a vector database and will be used for vector search in later steps. A well-developed vector database will properly index the data and optimize the retrieval process. The retrieval part starts by a user asking a question in the form of prompts to the same embedding model, which has generated the vector representation of the stored data, and gaining the vector embeddings of the question. The next step of the process is vector searching inside the vector databases The vector search is essentially computing similarity scores among vectors, and the database then identifies and retrieves the data segments that have the highest similarity scores (top K in most RAG systems) compared to the query vector. These retrieved segments are then converted back from their vector format to their original form, which is typically text of documents. In the genertation part, LLMs are involved to generate the final answers. The retrieved documents, along with the user’s question, are incorporated into a specifically chosen and designed prompt template. This template selection is based on the task type and aims to effectively merge the context with the question to form a coherent prompt. The selected LLM is provided with the prompt and generates the final answer.
3.2.2 RAG with Vector Databases
@@
In the prototype of the RAG paradigm, we use such an idea to overcome the hallucination problems of LLMs by providing domain-specific data with accurate instruction, where the vector database serves as an external knowledge base to warehouse domain-specific data, then LLMs can easily handle massive data owned by users.
@@
Even though LLMs such as ChatGPT now have user-specified GPTs for specific uses, with just prompt engineering, their knowledge bases are still limited to the training data provided by OpenAI. However, by using a vector database, users can pre-filter whatever they want it to look at, whereas that is difficult with prompt-engineering GPT assistants.
3.2.3 Expanding Horizons: Multi-modal Integration and Retrieval Innovations in RAG Systems
Numerous studies have been undertaken to enhance the performance and functionality of RAG systems.
Datasets
RAG was originally developed and evaluated on text-based question-answering (QA) tasks. There are numerous datasets and benchmarks available for evaluating the performance of various RAG systems, where different prompt templates need to be designed for different tasks. Therefore, we classified datasets based on the type of tasks and summarized them below.
  • Single-hop QA involves simple questions where the answer can be obtained through a single-step reasoning process, including TriviaQA Joshi and et al. (2017), Natural Questions Kwiatkowski and et al. (2019), WebQuestions Berant and et al. (2013). Multi-hop QA requires multiple intermediate reasoning steps to reach the final answer. For example, for a question like ”In which state did Obama go to high school?”, one needs to first infer which school Obama attended and then determine the location of that school to obtain the final answer. Datasets for multi-hop QA include HotpotQA Yang and et al. (2018), 2WikiMultiHopQA Ho and et al. (2020), MuSiQue Trivedi and et al. (2022), StrategyQA Geva and et al. (2021). In Multiple-choice QA, the model does not need to generate an answer independently but rather select the correct answer from provided options. Multiple-choice QA datasets include MMLU Hendrycks and et al. (2020), SQuAD Rajpurkar and et al. (2016). Open-domain QA involves the task of searching for answers to information-seeking queries within a large pool of knowledge sources. Relevant datasets for open-domain QA include MS MARCO Nguyen and et al. (2016), PopQA Mallen and et al. (2022), CommonsenseQA Talmor and et al. (2018).
  • Fact-checking requires the model to first select relevant sentences from a given document as evidence and then validate statements based on that evidence. Datasets include FEVER Thorne and et al. (2018), FM2 Eisenschlos and et al. (2021).
  • Embedding evaluation comprehensively assess all essential capabilities of text embeddings, including BEIR Thakur and et al. (2021), C-MTEB Xiao and et al. (2023).
Multimodality of RAG
RAG has now evolved to handle a wide range of data types by lending the power of multimodal models.
The impressive achievements of LLMs have inspired significant advancements in vision-and-language research. DALL-E from OpenAI introduced a Transformer-based approach for converting text to images, treating images as sequences of discrete tokens. Subsequent improvements in the text-to-image area Zhang et al. (2023) have been achieved through methods like model scaling, pretraining, and enhanced image quantization models. BLIP-2 Li et al. (2023) uses static image encoders with LLMs for efficient visual language pre-training, facilitating direct image-to-text transformations. Flamingo Alayrac et al. (2022) presented a visual language model for text generation, showcasing remarkable adaptability and leading performance across various vision-and-language tasks. CM3 Aghajanyan et al. (2022) trained a randomly masked model on a large HTML corpus, and showed that the model is capable of generating images and text. FROMAGe Koh et al. (2023) gains robust multimodal capabilities for few-shot learning solely from image-caption pairs, unlike other models that necessitate large-scale, interwoven image-text data from the websites.
To import speech data to RAG systems, Wav2Seq Wu et al. (2022) allows for efficient pre-training without the need for transcriptions, using techniques like k-means clustering and byte-pair encoding to generate pseudo subwords from speech. The Large Language and Speech Model (LLaSM) Shu et al. (2023) is an end-to-end trained, large multi-modal speech-language model equipped with cross-modal conversational skills and proficient in understanding and responding to combined speech-and-language directives. Videos are also made available to certain types of RAG systems. Vid2Seq Yang et al. (2023) enhances language models with specific temporal indicators for predicting event limits and textual descriptions in a single output sequence.
Retrieval Optimizations of RAG
To better harness the knowledge from various types of data, kNN-LMs Khandelwal et al. (2020) explores how incorporating nearest neighbor search into language models can enhance their ability to generalize by effectively leveraging memorized examples. EASE Nishikawa et al. (2022) is distinctive in its use of entities as a strong indicator of text semantics, providing rich training signals for sentence embedding. In-context RALM Ram et al. (2023) proves that with the language model architecture unchanged, simply appending grounding documents to the input will improve the performance. SURGE Kang et al. (2023) enhances dialogue generation by incorporating context-relevant sub-graphs from a knowledge graph. Another work that combines knowledge graphs and LLMs is RET-LLM Modarressi et al. (2023), which is designed to equip large language models with a general write-read memory unit. These studies have focused on retrieval granularity and data structuring levels, with coarse granularity providing more, but less precise, information. Conversely, structured text retrieval offers detailed information at the cost of efficiency.
To utilize both internal knowledge and external resources, SKR Yu et al. (2023) improves LLMs’ ability by enabling them to assess what they know and identify when to seek external information to answer questions more effectively. Selfmem Cheng et al. (2023) enhances retrieval-augmented text generation by creating an unbounded memory pool using the model’s own output and selecting the best output as memory for subsequent rounds. FLARE Jiang et al. (2023) uses predictions of upcoming sentences to anticipate future content and retrieve relevant documents for regenerating sentences, especially when they contain low-confidence tokens. Atlas Izacard et al. (2022) demonstrates impressive performance in tasks like question-answering and fact-checking, outperforming much larger models despite having fewer parameters. Many other works like these also aim to make RAG system more efficient and competent.
@@ 3.3 VecDB as a Reliable Memory of GPTs
The integration of Large Language Models (LLMs) such as the Generative Pre-trained Transformer (GPT) with vector databases (VecDB) heralds a transformative leap in the field of information retrieval and artificial intelligence. This subsection delves into the compelling synergy between VecDB as a persistent, reliable memory layer and the dynamic, contextually aware computation layer provided by LLMs. In recent works of vector databases that integrate with LLMs, Zhang and et al. (2023) showcases that vector databases’ capability of storing vectors for GPT’s memory.
Vector databases serve as a robust memory layer that addresses one of the intrinsic limitations of LLMs: the static nature of their knowledge. While LLMs excel in generating human-like text based on patterns learned during training, they cannot inherently update their knowledge base dynamically. VecDBs bridge this gap by offering a storage solution that can be continually updated with new information, ensuring that the LLM’s responses are informed by the most current and relevant data available.
@@
3.4 VecDB as a Cost-effective Semantic Cache
The combination of vector databases and LLM not only facilitates the profound application of LLM with RAG but also provides a new frontier for cost-effective end-to-end LLM applications:
Outrageous API Costs
LLM-based chatbots and agent systems heavily rely on LLM’s output from API vendors, considerable repeated or similar inquiries may lead to outrageous API costs.
API Bandwidth Limitations
Such chatbots and agent systems could also experience bursty inquiries workload that may drown the system’s bandwidth with explosive API calls coming within seconds, leading to the system’s outage and reconfiguration.
One of the primary benefits of integrating vector databases with LLMs is the significant reduction in data operational costs. GPT-Cache, as shown in Figure 2, for instance, stores responses to previously asked queries, and works as a cache before calling LLM APIs. This caching mechanism means that the system doesn’t really need to have API calls to wait for generated responses from scratch every time, reducing the number of costly API calls to the LLM. Moreover, this approach also speeds up the response time, enhancing user experience.
Figure 2: An overview of semantic cache for GPTs that utilizes vector database
Vector databases enhance the LLMs’ ability to retrieve and utilize relevant information by indexing vast amounts of previous Q&A data and mapping them into a vector space, instead of caches in computer systems that require exact hash-match, this allows more precise semantic matching of queries with existing knowledge and results in responses that are not only generated based on the LLM’s training data but also informed by the most relevant and recent information available in the vector database.
@@ 4.0.1 Are vector searches all you need?
Although vector databases offer a cost-efficient modality for information retrieval within LLM frameworks, their utility in traditional relational database operations remains limited. Specifically, vector-based search methodologies are often not optimized for operations such as post-query filtering, comprehensive full-text searches, and nuanced keyword search mechanisms that are fundamental to conventional database systems. This distinction underscores a potential gap in the functional alignment of vector databases with established data retrieval paradigms, necessitating the development of hybrid search algorithms that can seamlessly integrate vector search with traditional relational database capabilities.
@@ 4.0.2 VDBMS with multi-modality
While uni-modal data may lack the prospect of providing informative and more contextually appropriate search results, the multi-modal data and its hybrid processing also present a non-trivial challenge for vector databases’ storage and retrieval Wang and et al. (2023), since integrating and merging multi-modal data require efficient preprocessing with multi-modal encoders, multi-modal storage indexes, but also weight assignment and multi-modal data fusion.

### v2

@@ 2.2.3 A feasible solution of knowledge base for LLMs
Nevertheless, VecDBs are capable of solving the aforementioned problem by simply transforming data into vectors and leveraging their retrieval efficiency; thus, the emergence of VecDBs (as shown in Table 1) has greatly influenced machine learning systems and their multi-modal applications. An intuitive application using vector search is searching images with an image on search engines built on VecDBs Wang et al. (2021) like Google Image Search  [^2] https://cloud.google.com/vertex-ai/docs/vector-search/overview, which uses one input image to find similar images on the internet. While VecDBs boast their unique search capability and efficiency for unstructured data, they are just a back-end factor for manifold applications. To maximize their abilities, this leads us to an interesting question: since LLMs use embeddings to represent text as vectors, can developers combine VecDBs and LLMs to overcome the aforementioned challenges that are inherited in LLM applications? The answer is yes.
@@
As a kind of database encapsulated with vector search in vector data that represents real-world information within high dimensionalities, VecDBs are well-capable for retrieval applications Asai et al. (2023) incorporating LLMs because LLM applications are generally read-intensive, not requiring many write-related changes, especially data deletes Pan et al. (2023). On the other hand, VecDBs can efficiently manage and warehouse vector data required and generated by LLMs, thus providing a solid data cornerstone for both LLMs and their applications. While LLMs are often limited by domain knowledge that cannot be uploaded or distributed due to security and privacy concerns, domain knowledge is often unstructured data that can easily be embedded into VecDBs for efficient local retrieval and further integration with generative AIs. Moreover, the computing and storage resources of VecDBs are way cheaper than LLMs, since they do require costly GPUs and TPUs, thus achieving a cost-effective way of fast retrieval and durable (non-volatile) storage.
3 Synergizing LLMs with VecDBs: Enhancements and Innovations
Figure 1: A sample RAG framework that uses VecDBs.
With recent rapid development in both areas, applications that utilize the combination of the two are also growing rapidly, and the most fruitful one is Retrieval-Augmented Generation (RAG), where VecDBs play a crucial role in enabling cost-effective data storage and efficient retrieval.
3.1 VecDB as an External Knowledge Base: Retrieval-Augmented Generation (RAG)
@@ 3.1.1 Development of RAG Paradigm
As the release of ChatGPT casts a spotlight on LLMs to the public world, there is an increasing need to use AI chatbots as query and retrieval agents. But simply loading users’ private data as input to LLMs is found to be incompetent in real-world applications. LLMs have been constrained by their limited token counts and the high costs associated with training and fine-tuning for every alteration of data, especially when dealing with personalized or business-specific responses that require real-time data updates. To address such concerns and needs, retrieval-augmented generation (RAG) emerges as a novel solution that addresses the challenges faced by LLMs in integrating and processing large and dynamic data in external databases, where VecDBs offer a solution by acting as external memory for LLMs. They allow for the segmentation of private data by converting them into vectors and storing these vectors efficiently for a quick retrieval process. Integrating with VecDBs enables LLMs to access and synthesize enormous amounts of data without the need for constant re-training, thereby overcoming their inherent limitations.
@@
The concept of RAG is devised to be a paradigm, and a common workflow of RAG is illustrated in Figure 1. There are essentially three main parts to a full run of the system: data storage, retrieval, and generation.
@@
Data storage means establishing reliable external memory like VecDBs, and there are detailed recipes for this process Han et al. (2023). It starts with data preprocessing, during which the original data is collected, cleaned, integrated, transformed, normalized, and standardized. The processed data is chunked into smaller segments because of the contextual limitations of LLMs. These segments of data are then converted by an embedding model into vectors, the semantic representations of the data, which are stored in VecDBs and will be used for vector search in later steps. A well-developed VecDB will properly index the data and optimize the retrieval process. The retrieval part starts with a user asking a question in the form of prompts to the same embedding model, which has generated the vector representation of the stored data and gained the vector embeddings of the question. The next step of the process is vector searching and scoring inside the VecDBs, which essentially involves computing similarity scores among vectors, and the database then identifies and retrieves the data segments that have the highest similarity scores (top $K$ in most RAG systems) compared to the query vector. These retrieved segments are then converted back from their vector format to their original form, which is typically the text of documents. In the generation part, LLMs are involved in generating the final answers. The retrieved documents, along with the user’s question, are incorporated into a specifically chosen and designed prompt template. This template selection is based on the task type and aims to effectively merge the context with the question to form a coherent prompt. The selected LLM is provided with the prompt and generates the final answer.
3.1.2 RAG with VecDBs
@@
In the prototype of the RAG paradigm, we use such an idea to overcome the hallucination problems of LLMs by providing domain-specific data with accurate instructions. where VecDBs serve as an external knowledge base to warehouse domain-specific data, then LLMs can easily handle massive data owned by users.
@@
Even though LLMs such as ChatGPT now have user-specified GPTs for specific uses, with just prompt engineering, their knowledge bases are still limited to the training data provided by OpenAI. However, by using VecDBs, users can pre-filter whatever they want it to look at, whereas that is difficult with prompt-engineering GPT assistants.
Many research and industry examples justified RAG’s profound effect. Azure AI Search (formerly Cognitive Search) developed its vector search using Qdrant  [^3] https://learn.microsoft.com/en-us/azure/search/vector-search-overview. Another example is Pinecone’s Canopy  [^4] https://www.pinecone.io/blog/canopy-rag-framework/, an open-source framework that leverages Pinecone’s VecDB to build and host production-ready RAG systems. Companies like Spotify and Yahoo have adopted Vespa, an AI-driven online VecDB, and Yahoo uses it to help users directly chat with their mailbox, ask questions about their mailbox, and tell it to do things for them. This uses personalized data integrated with LLMs to form a RAG system  [^5] https://vespa.ai/. These examples demonstrate the strength of RAG combined with VecDBs to address the unique challenges faced by business enterprises in managing and extracting value from diverse datasets but also showcase how the aforementioned difficulties faced by LLMs can be solved via integration with VecDBs.
3.2 VecDB as a Cost-effective Semantic Cache
The combination of VecDBs and LLM not only facilitates the profound application of LLM with RAG but also provides a new frontier for cost-effective end-to-end LLM applications.
Outrageous API Costs
LLM-based chatbots and agent systems heavily rely on LLM’s output from API vendors; repeated or similar inquiries may lead to outrageous API costs.
API Bandwidth Limitations
Such chatbots and agent systems could also experience a bursty inquiry workload that may drown the system’s bandwidth with explosive API calls coming within seconds, leading to the system’s outage and reconfiguration.
One of the primary benefits of integrating VecDBs with LLMs is the significant reduction in data operational costs Sanca and Ailamaki (2023). GPT-Cache Bang (2023), a VecDB that serves as a semantic cache, as shown in Figure 2, for instance, stores responses to previously asked queries and works as a cache before calling LLM APIs. This caching mechanism means that the system doesn’t need to have API calls to wait for generated responses from scratch every time, reducing the number of costly API calls to the LLM. Moreover, this approach also speeds up the response time, enhancing the user experience.
VecDBs enhance the LLMs’ ability to retrieve and utilize relevant information by indexing vast amounts of previous Q&A data and mapping them into a vector space. Instead of caches in computer systems that require exact hash-match, this allows more precise semantic matching of queries with existing knowledge and results in responses that are not only generated based on the LLM’s training data but also informed by the most relevant and recent information available in the VecDBs.
@@ 3.3 VecDB as a Reliable Memory of GPTs
VecDBs serve as a robust memory layer Zhang et al. (2023b) that addresses one of the intrinsic limitations of LLMs: the static nature of their knowledge. While LLMs excel in generating human-like text based on patterns learned during training, they cannot inherently update their knowledge base dynamically and, thus, may lack the few-shot learning ability. VecDBs bridge this gap by offering a storage solution that can be continually updated with new information, ensuring that the LLM’s responses are informed by the most current and relevant data available.
Figure 2: An overview of semantic cache for GPTs that utilizes VecDBs.
@@
4 Expanding Horizons: RAG Advancements
4.1 Multimodality of RAG
RAG has now evolved to handle a wide range of data types by lending the power of multimodal models. The impressive achievements of LLMs have inspired significant advancements in vision-language research. DALL-E from OpenAI introduced a Transformer-based approach for converting text to images, treating images as sequences of discrete tokens. Subsequent improvements in the text-to-image area  Zhang et al. (2023a) have been achieved through methods like model scaling, pre-training, and enhanced image quantization models. BLIP-2  Li et al. (2023b) uses static image encoders with LLMs for efficient visual language pre-training, facilitating direct image-to-text transformations. Flamingo  Alayrac et al. (2022) presented a visual language model for text generation, showcasing remarkable adaptability and leading performance across various vision-and-language tasks. CM3  Aghajanyan et al. (2022) trained a randomly masked model on a large HTML corpus and showed that the model is capable of generating images and text. FROMAGe  Koh et al. (2023) gains robust multimodal capabilities for few-shot learning solely from image-caption pairs, unlike other models that necessitate large-scale, interwoven image-text data from the websites.
To import speech data to RAG systems, Wav2Seq  Wu et al. (2022) allows for efficient pre-training without the need for transcriptions, using techniques like k-means clustering and byte-pair encoding to generate pseudo-subwords from speech. The Large Language and Speech Model (LLaSM)  Shu et al. (2023) is an end-to-end trained, large multi-modal speech-language model equipped with cross-modal conversational skills and proficient in understanding and responding to combined speech-and-language directives. Videos are also made available for certain types of RAG systems. Vid2Seq  Yang et al. (2023) enhances language models with specific temporal indicators for predicting event limits and textual descriptions in a single output sequence.
4.2 Retrieval Optimizations of RAG
To better harness the knowledge from various types of data, kNN-LMs  Khandelwal et al. (2020) explore how incorporating nearest neighbor search into language models can enhance their ability to generalize by effectively leveraging memorized examples. EASE  Nishikawa et al. (2022) is distinctive in its use of entities as a strong indicator of text semantics, providing rich training signals for sentence embedding. In-context RALM  Ram et al. (2023) proves that with the language model architecture unchanged, simply appending grounding documents to the input will improve the performance. SURGE  Kang and Kwak (2023) enhances dialogue generation by incorporating context-relevant sub-graphs from a knowledge graph. Another work that combines knowledge graphs and LLMs is RET-LLM  Modarressi et al. (2023), which is designed to equip LLMs with a general write-read memory unit. These studies have focused on retrieval granularity and data structuring levels, with coarse granularity providing more, but less precise, information. Conversely, structured text retrieval offers detailed information at the cost of efficiency.
To utilize both internal knowledge and external resources, SKR  Yu et al. (2023) improves LLMs’ ability by enabling them to assess what they know and identify when to seek external information to answer questions more effectively. Selfmem  Xin et al. (2023) enhances retrieval-augmented text generation by creating an unbounded memory pool using the model’s output and selecting the best output as memory for subsequent rounds. FLARE  Jiang et al. (2023) uses predictions of upcoming sentences to anticipate future content and retrieve relevant documents for regenerating sentences, especially when they contain low-confidence tokens. Atlas  Izacard et al. (2022) demonstrates impressive performance in tasks like question-answering and fact-checking, outperforming much larger models despite having fewer parameters. Many other works like these also aim to make the RAG system more efficient and competent.
@@ 5.0.1 Are vector searches all you need?
Although VecDBs offer a cost-efficient modality for information retrieval within LLMs’ frameworks, their utility in traditional relational database operations remains limited. Specifically, vector-based search methodologies are still not well optimized for operations such as post-query filtering, comprehensive full-text searches, and keyword search mechanisms that are fundamental to conventional database systems and search engines. In the context of diverse applications for different users with multiple needs, vector search, or any single search approach, is not a one-size-fits-all solution. For example, for searching exact transaction data, users may have queries of very specific keywords to look up in a huge pool of data; giving them semantic approximate results totally would not make sense, resulting in a one-step forward yet two-step back. On the other hand, traditional DBMSs are well-developed for efficiently handling structured data within their transactional and analytical applications, while dedicated VecDBs are not yet.
This distinction underscores a potential gap in the functional alignment of VecDBs with previously well-established data retrieval paradigms, necessitating the development of hybrid search algorithms that can seamlessly integrate vector search with traditional relational data engine capabilities.
@@ 5.0.2 VDBMS with multi-modality
In real-world data representation and search, we may look for multi-modal data entries that are not restricted to a specific kind of structured, semi-structured, or unstructured data, but with any random combination of them, e.g., users can look for a knowledge graph from an interacted prompt of image and description. While searches using uni-modal data may lack the prospect of providing informative and more contextually appropriate search results, multi-modal data search and its hybrid processing also present a non-trivial challenge for VecDBs’ storage and retrieval Wang and et al. (2023), since integrating and merging multi-modal data require efficient preprocessing with multi-modal encoders, multi-modal storage indexes, but also ranking weight assignment and multi-modal data fusion.